# Analysis and experiments of the dissipative Twistcar: direction reversal and asymptotic approximations

### v1

@@ V Experiments
We designed a robot with a modular reconfigurable geometric structure and mass distribution, as shown in figure 6(a). We used a goBILDA kit for the chassis, a servo motor for the relative angle actuation, and an Arduino microcontroller for controlling the servo’s movement.
The experiments were conducted using three main robot configurations, as we show in table I. Those values were taken from the robot’s CAD model. The ‘Nominal Robot’ is the first and basic version of the robot, which was used for the parameter fitting experiments. Then, we modified the robot’s geometry so that a direction reversal would be possible by adding a block mass $m_{0}$ onto link 1. In the ‘Backward Configuration’, we extended the length of link 2 so that the forward speed of the vehicle is still negative but close to zero. The ‘Forward Configuration’ is the same configuration while adding the mass block, incorporating the block mass and moment of inertia into the properties of link 1.
@@ V-B Parameter fitting
Figure 7: Comparison between experimental results and simulation results with fitted viscous dissipation coefficient of $c=0.234[Kg/s]$. (a) Relative angle between the two links. (b) Vehicle’s body-frame velocity. (c) Orientation angle.
First, we look at a specific experiment with an input frequency of 14 radians per second and amplitude of $\pi/6$ radians, shown in figure 7. From figure 7(a), we see that the tracked relative angle is different from the expected input angle by its mean value, amplitude, and frequency. We define the tracked relative angle parameters in (36).
@@
In this experiment, we find the tracked input angle’s parameters: $\Phi_{\text{Mean}}=4.75^{\circ}$, $\Phi_{\text{Amp}}=23.75^{\circ}$ and $\Omega=12.36[rad/s]$, obtained by using the results’ extremum points. We conduct a numerical simulation for the tracked input, while hand-tuning the viscous dissipation coefficient. We add to figure 7 the simulation results with the best-fitted coefficient of $c=0.234[Kg/s]$. We see that using this model, we were able to capture the behavior of the robot’s velocity to a great extent, but were not able to match the actual drift rate of the orientation angle.
We note that in figure 7 we take a specific time window. Outside of this time window, the vehicle does not reach a steady-state and behaves unexpectedly. This behavior can be explained by an outside effect that the robot is sensitive to, and we did not account for while performing the experiment. In the following analysis, we look at each experiment and try to identify a steady-state, such as in figure 7(b), which we can analyze and compare to simulations.
@@
For each experiment, we measure the actual steering angle $\phi(t)$, shown in figure 8(b), which deviates significantly from the commanded input values. We use the actual steering angles as the simulations’ input, and use the results to compare with the experiments. In figure 8(a), we show the mean displacement per cycle of each experiment. By hand-tuning the viscous dissipation coefficient, we achieve the best alignment between the experiments and simulations with a value of $c=0.4[Kg/s]$.
Figure 8: Dissipation coefficients estimation using frequency-dependent results. (a) Displacement per cycle of different input frequency experiments overlaid with simulation results using $c=0.4[Kg/s]$. (b) Captured steering angle properties for different experiments, in dashed line we plot the expected steering angle properties.
From the comparison, we notice that no damping coefficient value matches the results at both high and low input frequencies. In addition, experiments conducted at high frequencies exhibit significant variance between different experiments under the same input. These results highlight the sensitivity of the robot, as previously discussed, and complicate the process of reliable parameter fitting.
@@ V-C Direction reversal experiment
Next, we obtain a direction reversal of the robot’s movement by placing a mass block on it. In figure 9, we present the experiment results as captured using the tracking system [^1] This paper has supplementary downloadable material available at https://drive.google.com/file/d/1BL6pWKHyYNvPq4p_D62BVZR0ibrW55cn/view?usp=sharing, provided by the authors. This includes a theoretical and experimental overview of the direction reversal phenomenon. This material is 67.6 MB in size.. Figure 9(a) contains the robot’s forward speed versus time, while figure 9(b) contains snapshots of the robot appearing at fixed time intervals of 6 seconds.
@@
Figure 10: Comparison between experimental results and simulation results with fitted viscous dissipation coefficients of $c_{\parallel}=0.095[Kg/s]$, $c_{\perp}=4[Kg/s]$. (a) Vehicle’s body-frame rolling velocity. (b) Vehicle’s body-frame skid velocity. (c) Orientation angle.
@@ VI-A Wheels’ skidding
In figure 10(a,c) we show the results of the same experiment as in figure 7, but now we also show the skid velocity $v_{\perp}(t)$ of the rear wheels in figure 7(b). The skid velocity of the vehicle is smaller by one order of magnitude than the forward speed, meaning it is small but not completely negligible.
To simulate this unconstrained model, we remove the nonholonomic no-skid constraints and incorporate a skid dissipation in the same way we incorporate rolling dissipation. We rewrite the skid velocity in terms of the Jacobian:
$\textbf{v}_{\perp,i}=\mathbb{J}_{\perp,i}(\textbf{q})\dot{\textbf{q}}$ (37)
Using the Jacobian forms in (4) and (37), we write the Rayleigh’s dissipation function, when $c$ is the viscous dissipation coefficient at the roll direction of each wheel and $c_{\perp}$ is the viscous dissipation coefficient at the skid direction.
@@
We hand-tune the viscous dissipation coefficients, and then choose the best-fitted coefficients $c_{\parallel}=0.095[Kg/s]$ and $c_{\perp}=4[Kg/s]$. The rolling dissipation coefficient obtained in this section is much smaller than the one that was obtained using the constrained model. This difference may explain the difference between the simulation and the experiments.
@@ VII Conclusions
The analysis involved perturbation expansion and was validated using numeric simulations. Using the analysis, we designed a modular robot demonstrating direction reversal. To determine the unknown viscous dissipation coefficient of the model, we conducted multiple experiments over a wide range of input frequencies, and also ran our simulation with input that was modified to match the actual measurements of joint angle in the experiment. By hand-tuning this coefficient, we found the best-fitting one as measured experimentally.
At the end of the work, we discussed some limitations of the model. First, one of our assumptions is no wheels’ skidding. However, experimental results indicate that skidding is present and not negligible. Second, we showed that the levelness of the ground’s surface can greatly influence the behavior of the robot.

### v2

@@ V Experiments
For the experiment, we designed a robot with a modular reconfigurable geometric structure and mass distribution, as shown in Figure 6(a). We used a goBILDA kit for the chassis, a servo motor for the relative angle actuation, and an Arduino microcontroller for controlling the servo’s movement.
The experiments were conducted using three main robot configurations, with physical parameter values as shown in Table I. Those values were taken from the robot’s CAD model. The ‘Nominal Robot’ configuration is the first and basic version of the robot, which was used for the parameter fitting experiments. Then, we modified the robot’s geometry so that a direction reversal would be possible by adding a block mass $m_{0}$ onto link 1. In the ‘Backward Configuration’, we extended the length of link 2 so that the forward speed of the vehicle is still negative but close to zero. The ‘Forward Configuration’ is the same configuration while adding the mass block, incorporating the block mass and moment of inertia into the properties of link 1.
@@ V-B Parameter fitting
Figure 7: Comparison between experimental results and simulation results with fitted viscous dissipation coefficient of $c=0.234[N\cdot s/m]$. (a) Relative angle between the two links. (b) Vehicle’s body-frame velocity. (c) Orientation angle.
First, we look at a specific experiment with an input frequency of 14 radians per second and amplitude of $\pi/6$ radians, shown in Figure 7. From Figure 7(a), we understand that the tracked relative angle $\phi(t)$ is different from the commanded input angle by its mean value, amplitude, and frequency. We define the tracked relative angle parameters in (38).
@@
In this experiment, we find the tracked input angle’s parameters: $\Phi_{\text{Mean}}=4.75^{\circ}$, $\Phi_{\text{Amp}}=23.75^{\circ}$ and $\Omega=12.36[rad/s]$, obtained by using the results’ extremum points. We conduct a numerical simulation under the tracked input while hand-tuning the viscous dissipation coefficient. We add to Figure 7 the simulation results with the best-fitted coefficient of $c=0.234[N\cdot s/m]$. Using this model, we were able to capture the behavior of the robot’s forward speed $v(t)$ to a great extent, but were not able to match the actual drift rate of the orientation angle $\theta(t)$.
We note that in Figure 7 we take a specific time window from the experimental measurements. Outside of this time window, the vehicle does not reach a steady-state motion and behaves irregularly. This behavior can be explained by external effects that the robot is sensitive to, and we did not account for while performing the experiment. In the following analysis, we look at each experiment and try to identify a steady-state, such as in Figure 7(b), which we can analyze and compare to simulations.
@@
For each experiment, we measure the actual steering angle $\phi(t)$, shown in Figure 8(b), which deviates significantly from the commanded input values. We use the actual steering angles as the simulations’ input, and use the results to compare with the experiments. In Figure 8(a), we show the mean displacement per cycle of each experiment. By hand-tuning the viscous dissipation coefficient, we achieve the best alignment between the experiments and simulations for $c=0.4[N\cdot s/m]$.
Figure 8: Dissipation coefficients estimation using frequency-dependent results. (a) Displacement per cycle of different input frequency experiments overlaid with simulation results using $c=0.4[N\cdot s/m]$. (b) Captured steering angle properties for different experiments, in dashed line we plot the expected steering angle properties.
From the comparison, we notice that no damping coefficient value matches the results at both high and low input frequencies. In addition, experiments conducted at high frequencies exhibit significant variance between different experiments under the same input. These results highlight the sensitivity of the robot and complicate the process of reliable parameter fitting.
@@ V-C Direction reversal experiment
Next, we obtain a direction reversal of the robot’s movement by placing a mass block on it. In Figure 9, we present the experiment results as captured using the tracking system. The results are also shown in a supplementary movie [^1] This paper has supplementary downloadable material available at https://drive.google.com/file/d/1BL6pWKHyYNvPq4p_D62BVZR0ibrW55cn/view?usp=sharing, provided by the authors. This includes a theoretical and experimental overview of the direction reversal phenomenon. This material is 67.6 MB in size.. Figure 9(a) contains the robot’s forward speed versus time, while Figure 9(b) contains snapshots of the robot appearing at fixed time intervals of 6 seconds.
@@
Figure 10: Comparison between experimental results and simulation results with fitted viscous dissipation coefficients of $c_{\parallel}=0.095[N\cdot s/m]$, $c_{\perp}=4[N\cdot s/m]$. (a) Vehicle’s body-frame rolling velocity. (b) Vehicle’s body-frame skid velocity. (c) Orientation angle.
@@ VI-A Wheels’ skidding
In Figure 10(a,c), we show the results of the same experiment as in Figure 7, and add in Figure 7(b) the measured skid velocity $v_{\perp}(t)$ of the rear wheels. The skid velocity of the vehicle is smaller by one order of magnitude than the forward speed, meaning it is small but not completely negligible.
In the following analysis, we incorporate wheels’ skid into the model by replacing the nonholonomic no-skid constraints with skid dissipation. We use the Jacobian forms in (3) and (6), and write the Rayleigh’s dissipation function when $c$ is the viscous dissipation coefficient at the roll direction of each wheel and $c_{\perp}$ is the viscous dissipation coefficient at the skid direction:
@@
We hand-tune the viscous dissipation coefficients, and choose the best-fitted coefficients $c_{\parallel}=0.095[N\cdot s/m]$ and $c_{\perp}=4[N\cdot s/m]$, as presented in Figure 10. The rolling dissipation coefficient obtained in this section is much smaller than the one that was obtained using the constrained model. This difference may explain the difference between the simulation and the experiments.
@@ VII Conclusions
The analysis involved perturbation expansion and was validated using numeric simulations. Based on the theoretical analysis, we designed a modular robot demonstrating direction reversal. To determine the unknown viscous dissipation coefficient of the model, we conducted multiple experiments over a wide range of input frequencies, and also ran our simulation with input that was modified to match the actual measurements of joint angle in the experiment. By hand-tuning this coefficient, we found the best-fitting one as measured experimentally.
At the end of the work, we discussed some limitations of the model, as well as the influence of some effects that have not been accounted in the model. First, we incorporated the effect of wheels’ skidding into the numerical simulations, which improved agreement with experimental measurements. Second, we showed that the levelness of the ground’s surface can greatly influence the behavior of the robot.